OpenStudy (anonymous):

Auditions for a play in New York were held with 120 men and women trying out for the lead singer position. If twice the number of men less three times the number of women is equal to 20, how many women auditioned? 44 40 76 80

OpenStudy (nurali):

m+w=120----(1) 3w-2m=20---(2) Multiply equation (1) by 2 and add 2m+2w=240 3w-2m=20 5w=260 w=260/5 w=52 ... 52 women put w in equation (1) m+w=120 ---> m+52=120 ----> m=68 ... 68 men

OpenStudy (nurali):

Welcome to Openstudy.